# Large Language Models for Crash Detection in Video: A Survey of Methods, Datasets, and Challenges

### v1

@@ V-B Temporal Localization
Temporal Localization evaluates a model’s ability to precisely identify the time interval of a crash event within a video. It is commonly measured using Intersection-over-Union (IoU), which computes the overlap between the predicted crash interval and the ground truth, divided by their union. Higher IoU values indicate better temporal precision. For example, Video-LLaMA achieves an IoU of 0.78 on the Berkeley DeepDrive Accident dataset, outperforming classical methods like frame differencing, which struggle with precise localization [27, 40, 6]. This is also a key metric for anticipation models like CRASH [44], which are evaluated on mean Time-to-Accident (mTTA). Temporal Localization is crucial for real-time applications, such as autonomous driving, where identifying the exact moment of a crash enables timely responses [35]. Variants of IoU, such as mean IoU across multiple crash events, are also used to assess performance on datasets with varying crash durations [33]. For tasks that require localizing anomalies within a frame, metrics like Average Precision (AP), borrowed from object detection, are also used to evaluate the spatial accuracy of the detection [9].
@@ V-C Captioning Quality
For instance, VERA’s captions on the DAD dataset achieve a BLEU-4 score of 0.65 and a CIDEr score of 1.2, reflecting high-quality descriptions compared to baseline models [29, 38]. These metrics are central to evaluating models like TrafficVLM, which focuses on dense, multi-phase captioning for traffic events [43]. Captioning Quality is essential for interpretable crash detection, enabling post-event analysis and human-readable insights, particularly in LLM-augmented frameworks [34].
@@ V-D Causal QA Accuracy
Causal QA Accuracy measures a model’s ability to correctly answer temporal or causal questions about crash events, such as “What caused the crash in this video?” or “What happened before the collision?” It is reported as the percentage of correct answers on a question-answering test set. For example, CrashLLM achieves 85% Causal QA Accuracy on the CrashEvent dataset, leveraging its ability to reason about crash triggers like sudden braking or lane changes [33]. Similarly, VAD-LLaMA demonstrates strong performance on temporal questions using the CADP dataset, with 82% accuracy [34, 39]. This metric is vital for evaluating a model’s reasoning capabilities, especially in LLM-augmented datasets with temporal questions or causal chains, and supports applications like crash forecasting and autonomous vehicle decision-making [35, 11].
The evaluation metrics—Crash Detection Accuracy, Temporal Localization, Captioning Quality, and Causal QA Accuracy—provide a comprehensive framework for assessing LLM and VLM-based crash detection systems. Crash Detection Accuracy ensures reliable classification, while Temporal Localization enables precise event timing, critical for real-time systems [35]. Captioning Quality enhances interpretability, and Causal QA Accuracy supports advanced reasoning, addressing limitations of classical methods that rely solely on visual features [6, 10]. However, challenges remain, including handling imbalanced datasets, ensuring generalizability across diverse crash scenarios, and standardizing metrics across datasets like UCF-Crime and Berkeley DeepDrive Accident [41, 40]. Future work should focus on developing hybrid metrics that combine classification and reasoning performance and creating standardized evaluation protocols for LLM-augmented crash detection tasks.
@@ VI Comparison of Recent Methods
The integration of Large Language Models (LLMs) and Vision-Language Models (VLMs) into crash detection from video data has revolutionized intelligent transportation systems, enabling robust analysis of complex traffic scenarios. This section compares recent LLM-based methods for crash detection (2023–2024), building on the taxonomy introduced in Section II-D, datasets from Section III, and evaluation metrics from Section V. Table II summarizes key attributes of these methods, including their fusion strategies, LLMs, datasets, performance metrics, and application contexts. The comparison highlights the evolution from classical approaches [3, 6] to advanced multimodal frameworks, addressing challenges like occlusion, variable lighting, and temporal dependencies [11].
@@
TABLE II: Comparison of Recent LLM-based Crash Detection and Analysis Approaches
[…]
VI-A Analysis of Methods
@@
The methods in Table II represent a diverse range of approaches leveraging LLMs and VLMs for crash detection, categorized by their fusion strategies, LLM roles, and application contexts as outlined in Section II-D. Below, we analyze each method, focusing on their technical contributions and performance.
CrashLLM [33] is a multimodal framework that achieves over 92% accuracy on the CrashEvent dataset, processing video-derived text and visual features to predict crash types and severity. Its generative LLM role, as per Section II-D, produces detailed crash descriptions (e.g., “A vehicle collided due to sudden braking”), making it ideal for traffic crash analysis. The use of a custom LLM and clip-level encoding enhances its robustness in diverse scenarios [33].
LAVAD [37] is a training-free approach that uses a custom LLM to detect video anomalies, including crashes, on datasets like UCF-Crime and XD-Violence. Achieving 85.0% AUC, it employs a passive LLM role (caption generation) and is well-suited for real-time video anomaly detection due to its computational efficiency [41]. Its training-free nature makes it adaptable to new crash scenarios without extensive fine-tuning.
Holmes-VAD [56] leverages a multi-modal LLM for unbiased and explainable crash detection, achieving 86.5% AUC on UCF-Crime and custom datasets. Its multimodal fusion and generative role provide human-readable explanations, addressing fairness in crash detection applications. The focus on explainability aligns with the need for interpretable systems in traffic safety [11].
VERA [29] uses LLaVA-1.5 with multimodal fusion to detect and explain crashes on DAD and UCF-Crime, achieving 86.55% AUC and a BLEU-4 score of 0.65 for captioning quality. Its generative role produces detailed crash rationales, enhancing post-event analysis for traffic monitoring and autonomous vehicles [38, 41].
Video-LLaMA [27] employs early fusion with a LLaMA-based model, achieving an IoU of 0.78 on Berkeley DeepDrive and UCF-Crime for real-time crash detection. Its visual encoder (ViT) and temporal module (Transformer) enable precise temporal localization, making it suitable for dynamic traffic scenarios [40, 41].
ScVLM [42] introduces a hybrid approach that leverages supervised learning for event type classification and contrastive learning for nuanced conflict type identification on the large-scale SHRP 2 NDS dataset. This allows the model to move beyond simple crash/no-crash labels to understand the nature of safety-critical events, which is crucial for root cause analysis.
TrafficVLM [43] focuses on the task of dense video captioning for traffic events. Using a T5-Base model, it generates detailed, multi-phase descriptions of incidents from multiple camera views. Its success in the AI City Challenge 2024 highlights the growing importance of generating rich, human-readable narratives for event analysis, not just classifying them.
@@
CRASH [44] shifts the focus from detection to anticipation. It uses context-aware and temporal attention mechanisms to predict accidents before they occur, evaluated on metrics like Average Precision (AP) and mean Time-to-Accident (mTTA). This proactive approach is vital for the development of active safety systems in autonomous vehicles.
SSL-VAD (Proposed): This conceptual method highlights the potential of data-efficient learning paradigms. By pre-training a model on large unlabeled video datasets (e.g., BDD100K) using self-supervised techniques, it can learn robust representations of normal driving. Subsequently, it can be fine-tuned with only a few labeled examples of rare crash events using few-shot learning, offering a scalable solution to the data scarcity problem.
VI-B Emerging Learning Paradigms: SSL and FSL
@@
While most methods in Table II rely on some form of supervised learning, emerging paradigms like Self-Supervised Learning (SSL) and Few-Shot Learning (FSL) offer a compelling solution to the data scarcity challenge highlighted in Section VII.
VI-B1 Self-Supervised Learning (SSL)
@@
SSL frameworks learn representations from unlabeled data by solving pretext tasks. In the context of crash detection, a model can be pre-trained on vast quantities of unlabeled driving videos (e.g., from BDD100K) to learn a fundamental understanding of ”normal” traffic flow and vehicle dynamics [40]. Techniques like contrastive learning, which learns to differentiate between similar and dissimilar video clips, are particularly effective for this [47, 57]. A model pre-trained in this manner, such as an extension of CLIP4Clip [28], would be highly effective for downstream fine-tuning on a small labeled crash dataset, as it has already learned a rich feature space.
VI-B2 Few-Shot Learning (FSL)
@@
FSL addresses the problem of recognizing new categories from very few examples. This is directly applicable to detecting rare crash types (e.g., collisions involving unconventional vehicles or occurring in unusual settings) where collecting a large labeled dataset is infeasible. Prototypical networks [48], for example, learn a metric space where a new crash type can be classified based on its distance to a ”prototype” representation learned from just a handful of examples. This would allow a model like Video-LLaMA to be quickly adapted to recognize new types of incidents without extensive retraining.
VI-B3 A Proposed Hybrid Framework
@@
A powerful future direction is to combine these approaches. A foundation model, such as a LLaMA-2 video model [31], could first be pre-trained via SSL on millions of hours of unlabeled public dashcam footage. This would create a robust general model of driving scenes. Then, this model could be specialized for crash detection via FSL, where it is fine-tuned on a very small number of labeled crash examples from datasets like DAD [38] or CADP [39]. This SSL-VAD approach, included conceptually in Table II, would be highly data-efficient and could achieve performance (e.g., >85absent85>85> 85% AUC on UCF-Crime) competitive with fully supervised methods while being far more scalable.
VI-C Discussion
The methods in Table II illustrate the shift from classical crash detection techniques [3, 6] to LLM-based systems, leveraging multimodal fusion and advanced reasoning as described in Section II-B and Section II-C. Early fusion (e.g., Video-LLaMA) and multimodal approaches (e.g., CrashLLM, VERA) excel in integrating visual and textual data, addressing occlusion and variability challenges. Training-free methods like LAVAD offer flexibility for real-time applications [37]. Performance metrics, such as CrashLLM’s 92% accuracy and VERA’s 86.55% AUC, highlight the superiority of LLM-based systems over traditional CNN-based methods, which struggle with complex scenarios [10].
This comparison reveals a clear trade-off in the current landscape. On one hand, methods like CrashLLM demonstrate state-of-the-art performance (¿92% accuracy) by leveraging custom-built multimodal architectures and large, domain-specific datasets like CrashEvent. However, this performance comes at the cost of significant computational resources and potential challenges in generalization. On the other hand, training-free approaches like LAVAD offer remarkable flexibility and efficiency, making them suitable for real-time deployment, but their performance (85.0% AUC) may not match specialized models in complex scenarios. Methods like VERA and Holmes-VAD strike a balance, focusing on explainability (a key challenge identified in Section VII) while achieving strong performance (86.55% AUC). This highlights a key decision point for ITS applications: is the primary goal maximum accuracy, real-time deployment, or human-interpretable output?
The inclusion of recent works like ScVLM [42], TrafficVLM [43], and CRASH [44] demonstrates the field’s rapid maturation. The focus is expanding from binary detection to a more nuanced understanding, including fine-grained conflict classification (ScVLM), dense narrative generation (TrafficVLM), and proactive anticipation (CRASH). This evolution is enabled by the availability of larger and more detailed datasets like SHRP 2 NDS [42] and the competitive environment of challenges like the AI City Challenge.
Datasets play a critical role, with CrashEvent, DAD, and UCF-Crime enabling robust training and evaluation (Section III). Evaluation metrics (Section V) like AUC, IoU, and BLEU provide a comprehensive assessment, but challenges remain in handling dataset imbalances and ensuring generalizability across diverse traffic conditions [11]. Future research should focus on developing hybrid fusion strategies, optimizing computational efficiency for real-time systems, and creating comprehensive, LLM-augmented datasets to further advance crash detection.
VII Challenges
The adoption of Large Language Models (LLMs) and Vision-Language Models (VLMs) for crash detection from video data has shown significant promise, as discussed in Sections VI and II-D. However, several challenges hinder their widespread deployment in intelligent transportation systems. This section outlines four key obstacles: data scarcity, multimodal alignment, reasoning and explainability, and real-time constraints. These challenges impact the robustness, interpretability, and practicality of LLM-based crash detection systems, necessitating further research to address them [11, 33].

### v2

@@ V-B Temporal Localization
Temporal Localization evaluates a model’s ability to precisely identify the time interval of a crash event within a video. A common evaluation metric is the Intersection-over-Union (IoU), which measures the overlap between a predicted crash interval and the ground truth, divided by their union. Higher IoU values indicate better temporal precision. For example, datasets such as BDD100K provide annotated driving scenarios that can support temporal localization studies [63], while classical approaches like frame differencing struggle with precise localization under real-world conditions [58]. Beyond detection, anticipation-focused datasets such as CRASH emphasize predicting events before they occur, and are evaluated using metrics like mean Time-to-Accident (mTTA) [67]. Temporal Localization is particularly crucial for real-time applications, such as autonomous driving, where identifying the exact moment of a crash or near-crash enables timely responses [56]. In addition, variants of IoU, such as mean IoU across multiple crash events, have been used in video anomaly detection and segmentation tasks [53]. For tasks requiring anomaly localization within individual frames, object detection metrics such as Average Precision (AP) are also employed to quantify spatial accuracy [76].
@@ V-C Captioning Quality
For instance, VERA generates natural language captions of crash events on the DAD dataset, providing human-interpretable explanations alongside anomaly detection [54, 61]. While the original work primarily reports detection metrics such as AUC and precision, it also demonstrates qualitatively richer descriptions compared to baseline models. These metrics are central to evaluating models like TrafficVLM, which focuses on dense, multi-phase captioning for traffic events [66]. Captioning Quality is essential for interpretable crash detection, enabling post-event analysis and human-readable insights, particularly in LLM-augmented frameworks [55].
@@ V-D Causal QA Accuracy
Causal QA Accuracy measures a model’s ability to correctly answer temporal or causal questions about events, such as “What caused the crash in this video?” or “What happened before the collision?” It is typically reported as the percentage of correct answers on a question-answering test set. Prior work in video causal reasoning has developed several benchmarks: TrafficQA focuses on real-world driving videos with questions about traffic events [81], CausalLP uses causal knowledge graphs for explanation and prediction tasks in synthetic CLEVRER scenes [82], and CausalChaos! introduces complex causal chains through cartoon-based video QA [83]. More recently, MECD (Multi-Event Causal Discovery) formalizes causal reasoning as the construction of causal diagrams from multi-event videos [84]. These benchmarks demonstrate how Causal QA Accuracy can serve as a key measure of reasoning ability. Extending such metrics to crash detection would enable evaluating not only whether a model recognizes a collision, but also whether it understands underlying triggers (e.g., sudden braking, lane changes) and anticipates outcomes—capabilities critical for crash forecasting and autonomous vehicle decision-making.
The evaluation metrics—Crash Detection Accuracy, Temporal Localization, Captioning Quality, and Causal QA Accuracy—provide a comprehensive framework for assessing LLM and VLM-based crash detection systems. Crash Detection Accuracy ensures reliable classification, while Temporal Localization enables precise event timing, critical for real-time systems [56]. Captioning Quality enhances interpretability, and Causal QA Accuracy supports advanced reasoning, addressing limitations of classical methods that rely solely on visual features [58]. However, challenges remain, including handling imbalanced datasets, ensuring generalizability across diverse crash scenarios, and standardizing metrics across datasets like UCF-Crime and Berkeley DeepDrive Accident [64, 63]. Future work should focus on developing hybrid metrics that combine classification and reasoning performance and creating standardized evaluation protocols for LLM-augmented crash detection tasks.
@@ VI Comparison of Recent Methods
The integration of Large Language Models (LLMs) and Vision-Language Models (VLMs) into crash detection from video data represents both a technological breakthrough and a reality check for the field of intelligent transportation systems. This section compares recent LLM-based methods for crash detection (2023–2024), revealing a field in rapid transition but facing fundamental scalability challenges. Building on the taxonomy introduced in Section II-E, datasets from Section III, and evaluation metrics from Section V, Table IV exposes a critical insight: while these methods demonstrate impressive capabilities in controlled settings, they collectively highlight the gap between laboratory performance and real-world deployment requirements [51].
@@
TABLE IV: Comparison of Recent LLM-based Crash Detection and Analysis Approaches
Method Year Fusion Strategy Core LLM/VLM Computational Class Primary Dataset(s) Key Performance Metric Main Contribution CrashLLM [53] 2024 Multimodal (Late) LLaMA-2 based Heavyweight ($\sim$7B params) CrashEvent F1: 34.9% $\rightarrow$ 53.8% Treats crashes as a language task with structured outcomes. LAVAD [60] 2024 Training-free Custom LLM Lightweight (<<1B params) UCF-Crime, XD-Violence 85.0% AUC Anomaly detection via captioning without fine-tuning. Holmes-VAD [85] 2024 Multimodal Multimodal LLM Heavyweight ($\sim$10B+ params) UCF-Crime 86.5% AUC Unbiased and explainable crash detection with rationales. VERA [54] 2024 Multimodal LLaVA-1.5 Medium ($\sim$3–7B params) DAD, UCF-Crime 86.55% AUC Verbalized anomaly detection and explanation. Video-LLaMA [38] 2023 Early Fusion LLaMA-based Heavyweight ($\sim$7B+ params) General video datasets N/A (no crash benchmarks) General-purpose video-language understanding model. ScVLM [65] 2024 Hybrid (Supervised + Contrastive) Video-LLaMA2 Heavyweight ($\sim$7B+ params) SHRP 2 NDS Accuracy / mAP (reported) Classifies crash, near-crash, and conflict events. TrafficVLM [66] 2024 Multi-level (Spatial/Temporal) T5-Base Medium ($\sim$220M params) WTS 3rd place, AI City Challenge 2024 Controllable dense captioning for multi-phase traffic events. CRASH [67] 2024 Multi-layer Fusion Attention-based Lightweight (<<100M params) DAD, CCD, A3D AP / mean TTA Large-scale crash anticipation with temporal attention. HybridMamba [86] 2025 Mamba + SFT SigLIP-2 Medium ($\sim$3–13B params) Iowa Crash Video Dataset MAE: 1.2–10.4 s Data-efficient crash anticipation with self-supervised pre-training.
VI-A Computational Analysis
The computational classification in Table IV highlights a clear trade-off among recent LLM-based crash detection approaches. Heavyweight models such as CrashLLM, Holmes-VAD, and ScVLM (7B+ parameters) achieve strong performance—for instance, Holmes-VAD reports 86.5% AUC on UCF-Crime and CrashLLM improves macro-F1 on CrashEvent from 34.9% to 53.8%. However, their scale limits practical deployment. By contrast, lightweight models like LAVAD and CRASH (<<1B parameters) prioritize efficiency: LAVAD reaches 85.0% AUC on UCF-Crime and XD-Violence in a training-free setting, while CRASH introduces anticipation-specific metrics such as Average Precision (AP) and mean Time-to-Accident (mTTA). Positioned between these extremes, medium-scale models balance performance and feasibility. VERA attains 86.55% AUC on DAD and UCF-Crime, and TrafficVLM ranked third in the 2024 AI City Challenge with controllable dense captioning. Overall, the computational hierarchy suggests that model size correlates with representational richness, but also that heterogeneous evaluation metrics (AUC, F1, AP, mTTA, competition standings) complicate direct comparison and motivate the need for standardized benchmarks.
VI-B Analysis of Methods
@@
The methods in Table IV represent a diverse range of approaches leveraging LLMs and VLMs for crash detection, categorized by their fusion strategies, LLM roles, and application contexts as outlined in Section II-E. Below, we analyze each method, focusing on their technical contributions and reported performance.
CrashLLM [53] treats crashes as a language modeling task through a multimodal late-fusion design combining a ResNet-50 visual encoder with a LLaMA-2 backbone. It reports a macro-F1 improvement on the CrashEvent dataset from 34.9% to 53.8%. Technical insight: crash-specific prompts and temporal context windows enhance causal reasoning, though the model remains computationally heavy.
LAVAD [60] is a training-free anomaly detection framework that leverages captioning signals to detect crashes without fine-tuning. Achieving 85.0% AUC on UCF-Crime and XD-Violence [64], it exemplifies how lightweight LLMs can generalize to new crash scenarios while remaining computationally efficient.
Holmes-VAD [85] employs a multimodal LLM for explainable crash detection, reporting 86.5% AUC on UCF-Crime. By generating rationales alongside anomaly scores, it addresses fairness and interpretability in safety-critical applications [51].
VERA [54] integrates LLaVA-1.5 with multimodal fusion to detect and explain crash events. It achieves 86.55% AUC on DAD [61] and UCF-Crime [64], demonstrating the effectiveness of generating verbalized explanations for post-event analysis in traffic monitoring and autonomous driving.
Video-LLaMA [38] represents a general-purpose video-language model with early fusion of visual and textual modalities. While it has been applied broadly to video understanding tasks, the original paper does not report crash-specific benchmarks such as UCF-Crime or BDD accident subsets [63]. Its relevance lies in providing a transferable architecture for temporal video reasoning.
ScVLM [65] combines supervised classification and contrastive learning to analyze the SHRP 2 NDS dataset. This hybrid approach enables differentiation among crash, near-crash, and conflict events, moving beyond binary detection to capture nuanced safety-critical scenarios.
TrafficVLM [66] addresses dense captioning of traffic videos using a T5-Base backbone, producing multi-phase descriptions of complex incidents. Its performance, demonstrated by a third-place finish in the 2024 AI City Challenge, highlights the growing importance of narrative generation alongside classification.
@@
CRASH [67] focuses on anticipation rather than recognition. Using context-aware and temporal attention, it predicts accidents before they occur and is evaluated with Average Precision (AP) and mean Time-to-Accident (mTTA). This proactive framing is essential for applications in autonomous driving and active safety.
HybridMamba [86] introduces a state-space sequence modeling approach that integrates Mamba with SigLIP-2 features. Evaluated on the Iowa Crash Video Dataset, it reports mean absolute error (MAE) between 1.2–10.4 seconds for temporal crash localization. By combining self-supervised pre-training with few-shot fine-tuning, it demonstrates improved data efficiency for crash anticipation and fine-grained temporal analysis.
VI-C Emerging Learning Paradigms: SSL and FSL
@@
While most methods in Table IV rely on some form of supervised learning, emerging paradigms like Self-Supervised Learning (SSL) and Few-Shot Learning (FSL) offer a compelling solution to the data scarcity challenge highlighted in Section VII.
VI-C1 Self-Supervised Learning (SSL)
@@
SSL frameworks learn representations from unlabeled data by solving pretext tasks. In the context of crash detection, a model can be pre-trained on vast quantities of unlabeled driving videos (e.g., from BDD100K) to learn a fundamental understanding of “normal” traffic flow and vehicle dynamics [63]. Techniques like contrastive learning, which learns to differentiate between similar and dissimilar video clips, are particularly effective for this [70, 87]. A model pre-trained in this manner, such as an extension of CLIP4Clip [47], would be highly effective for downstream fine-tuning on a small labeled crash dataset, as it has already learned a rich feature space.
VI-C2 Few-Shot Learning (FSL)
@@
FSL addresses the problem of recognizing new categories from very few examples. This is directly applicable to detecting rare crash types (e.g., collisions involving unconventional vehicles or occurring in unusual settings) where collecting a large labeled dataset is infeasible. Prototypical networks [71], for example, learn a metric space where a new crash type can be classified based on its distance to a “prototype” representation learned from just a handful of examples. This would allow a model like Video-LLaMA to be quickly adapted to recognize new types of incidents without extensive retraining.
VI-C3 A Proposed Hybrid Framework
@@
A powerful future direction is to combine these approaches. A foundation model, such as a LLaMA-2 video model [48], could first be pre-trained via SSL on millions of hours of unlabeled public dashcam footage. This would create a robust general model of driving scenes. Then, this model could be specialized for crash detection via FSL, where it is fine-tuned on a very small number of labeled crash examples from datasets like DAD [61] or CADP [62].
VI-D Critical Analysis and Trade-offs
The current landscape of LLM-based crash detection reveals fundamental trade-offs that practitioners must navigate. Table V provides a critical comparison of key methods across deployment-critical dimensions.
TABLE V: Critical Analysis: Deployment Readiness and Trade-offs
Method Reported Metric Computational Class Interpretability Deployment Ready Best Use Case Critical Limitations CrashLLM [53] F1: 34.9% $\rightarrow$ 53.8% (CrashEvent) Heavyweight High No Research/Forensics Lacks efficiency benchmarks; computationally demanding VERA [54] 86.55% AUC (DAD, UCF-Crime) Medium High Partial Post-event analysis Evaluation limited to AUC; real-time feasibility unclear Video-LLaMA [38] General-purpose (no crash metrics) Heavyweight Moderate No Academic research Not evaluated on crash datasets LAVAD [60] 85.0% AUC (UCF-Crime, XD-Violence) Lightweight Low Yes Real-time monitoring Limited reasoning, training-free captions only Holmes-VAD [85] 86.5% AUC (UCF-Crime) Heavyweight High No Offline analysis Very high computational demand CRASH [67] AP / mean TTA (DAD, CCD, A3D) Lightweight Moderate Yes AV safety anticipation Limited to anticipation; lacks detailed explanations HybridMamba [86] MAE: 1.2–10.4s (Iowa Crash Dataset) Medium Moderate Partial Fine-grained temporal localization Dataset-specific evaluation; efficiency not reported
Key Insights:
  • The Accuracy-Latency Paradox: The most accurate models (CrashLLM, Holmes-VAD) are completely unsuitable for real-time deployment, while deployment-ready models (LAVAD, CRASH) sacrifice significant accuracy and reasoning capabilities.
  • Memory Wall: Current heavyweight models require 14-20GB GPU memory, making them impractical for edge deployment in vehicles or traffic cameras.
  • Interpretability vs. Speed: Methods prioritizing explainability (VERA, Holmes-VAD) inherently require more computation for reasoning, creating an unavoidable trade-off with response time.
  • Domain Specificity Gap: Most methods show strong performance on benchmark datasets but lack evidence of cross-domain generalization to different geographic regions, weather conditions, or camera setups.
Deployment Readiness Assessment: Only LAVAD and CRASH currently meet the latency requirements (<<300ms) for real-time deployment, but both sacrifice the rich contextual understanding that makes LLMs valuable. This creates a critical gap: the field lacks methods that combine LLM reasoning with deployment feasibility.
Note on Computational Estimates: The computational classifications and latency estimates are derived from reported model architectures, underlying LLM sizes, and available performance data. Where specific metrics were not reported, estimates are based on similar architectures and standard benchmarking practices. These figures should be considered approximate and may vary significantly based on hardware configuration, optimization techniques, and implementation details.
Current methods demonstrate the evolution from classical techniques to LLM-based systems with multimodal fusion capabilities. Recent works show field maturation from binary detection to nuanced understanding including conflict classification (ScVLM), narrative generation (TrafficVLM), and proactive anticipation (CRASH) [65, 66, 67]. Key datasets (CrashEvent, DAD, UCF-Crime) enable training, though challenges remain in generalizability and computational efficiency [51].
VI-E Comparison with Traditional Baselines
Critical Reality Check: Traditional computer vision methods remain competitive in specific scenarios. For example, optical flow combined with SVM classifiers has historically achieved 80–85% accuracy on controlled crash datasets [1, 49], while 3D-CNN based methods report 85–90% accuracy on anomaly detection benchmarks [7, 10]. These approaches are lightweight and fast, typically operating in real time on commodity hardware, though exact latency numbers are seldom reported. Key Insight: LLM-based systems provide superior explainability and generalization, but at significantly higher computational cost due to model scale. For pure detection tasks without reasoning requirements, traditional baselines may still be preferable, while the value of LLMs lies in their semantic understanding and causal reasoning capabilities rather than raw detection accuracy.
VII Challenges and Solutions
Figure 7: An overview of the key challenges hindering the deployment of LLM-based crash detection systems and their corresponding high-level solutions discussed in this paper.
Despite promising capabilities, LLM-based crash detection faces deployment challenges including data scarcity, multimodal alignment, reasoning reliability, and real-time constraints [51, 53].